OpenStudy (anonymous):

Mikko manufactures drinking glasses in the shape of cylinders. He wants all of the glasses to have the same fixed volume. This implies that the height of a given glass will vary inversely with the square of its diameter. One of his glasses has a diameter of 3 inches and a height of 4 inches. If another glass has a height of 9 inches, what should its diameter be?

OpenStudy (anonymous):

easy as pie

OpenStudy (anonymous):

I can show you in my virtual classroom: http://www.twiddla.com/562616

OpenStudy (anonymous):

diameter will be 2 in

OpenStudy (anonymous):

You didn't show your work Harkirat

OpenStudy (anonymous):

ok here it is volume of first glass will be π*(3/2)^2 * 4 = π * 9/4 * 4 = 9π let radius of second glass be r its volume = π* r^2 *9 =9πr^2 since both have to have the same volume 9π = 9πr^2 ==> 9π/9π = r^2 = > 1 = r^2 so r = 1 diameter = 2r = 2*1 = 2

OpenStudy (darthsid):

Give a man a fish, he doesn't go hungry for a day. Teach him to fish, he never goes hungry. Or something like that. The height of the glass varies inversely with the square of the diameter. This means: \[height = (constant) \times \frac{1}{diameter^2}\] This means, \[height \times diameter^2\] = constant. Just make sure the products are the same for both, and you will be able to find your answer.